# WaterSeeker: Pioneering Efficient Detection of Watermarked Segments in Large Documents

### v3

@@ 3.1 Definition
Here, ‘has_watermark’ is a boolean value indicating the presence of a watermark, and ‘indices’ is a list of pairs $(s^{\prime},e^{\prime})$ representing the start and end indices of detected watermarked segments.
3.2 Evaluation
@@
$\text{IoU}=\cfrac{L_{\text{intersection}}}{L_{\text{union}}}>\theta.$
Based on this criteria, we will report the following metrics of the binary classification result: False Positive Rate (FPR), False Negative Rate (FNR), and F1 Score. Moreover, we will include the average IoU between the detected segments and ground truth segments as a supplementary metric, further demonstrating the accuracy of the localization.
4 Baseline Methods
4.1 Full-text Detection
@@
4.2 WinMax
WinMax (Kirchenbauer et al., 2024) involves iterating through all possible window sizes, and for each window size, the entire text is traversed to calculate the z-score for each local window, taking the maximum z-score and comparing it against a specified threshold. In practical applications, a minimum window size $W_{min}$ and a maximum window size $W_{max}$ are typically set to control the number of iterations during the traversal. The detection process can be described by the following formula:
$z_{\text{win-max}}=\max_{w\in[W_{min},W_{max}]}\left(\max_{i}\left(\frac{|s|_{G,i}-\gamma w}{\sqrt{\gamma(1-\gamma)w}}\right)\right),$
where $w$ is the length of the local window, and $|s|_{G,i}$ represents the count of green tokens within the $i$-th local window of length $w$. The time complexity of WinMax is $O((W_{max}-W_{min})\times N)$, where $N$ is the total length of the text. Due to the uncertain length of the watermarked segment inserted into the document, if there is a large difference between $W_{max}$ and $W_{min}$, the worst-case complexity can reach $O(N^{2})$.
4.3 Fix-Length Sliding Window
Fix-Length Sliding Window (FLSW) is a self-constructed method that uses a fixed-length window to traverse the text, calculating statistics within each local window. The text is flagged as watermarked if any statistic score exceeds the threshold. For localization, the method records each pair of indices $(s,e)$ meeting the condition and uses a contiguous segment merging method with tolerance. While straightforward, FLSW struggles with varying watermark lengths due to its fixed-length nature, leading to dilution effects. Section 5.1 analyzes this effect using KGW (Kirchenbauer et al., 2023) and Aar (Aaronson and Kirchner, 2022) as examples, providing theoretical support for WaterSeeker.
5 Proposed Method: WaterSeeker
5.1 Theoretical Basis: Gold Index is the Best
This section provides the theoretical foundation of WaterSeeker, showing that using actual start and end indices (gold index) for watermark detection achieves the highest expected detection rate. We use KGW and Aar as examples, since most watermarking algorithms are slight modifications of these two methods. Thus, analyzing these approaches offers broad applicability.
For the KGW method, assuming $\gamma_{1}>\gamma$ is the proportion of green tokens in the watermarked part. Let’s analyze the effect of window size $W$ on this statistic: (1) When $W<L$ (window size smaller than watermark length):
$E[z_{W}]=\frac{W\gamma_{1}-\gamma W}{\sqrt{\gamma(1-\gamma)W}}=\sqrt{W}\cdot\frac{\gamma_{1}-\gamma}{\sqrt{\gamma(1-\gamma)}}.$
(2) When $W>L$ (window size larger than watermark length):
$E[z_{W}]=\frac{L\gamma_{1}+(W-L)\gamma-\gamma W}{\sqrt{\gamma(1-\gamma)W}}=\frac{L(\gamma_{1}-\gamma)}{\sqrt{\gamma(1-\gamma)W}}.$
From this, we can conclude that when $W=L$, the z-score reaches its maximum.
During detection, we aim for a higher z-score for positive cases while setting an appropriate threshold to balance the false positive rate. Next, we will analyze the constraints on the z-threshold $z^{*}$ when the false positive rate within the specified window is set to be lower than a target value $\alpha$.
We start with the binomial distribution $B(W,\gamma)$ that describes the number of green tokens in a window of size $W$. For large sample sizes (typically $W>50$), we can approximate this as $|s|_{G}\sim N(W\gamma,W\gamma(1-\gamma))$. Therefore, we have $z\sim N(0,1)$. The false positive rate $\alpha$ represents the area in the right tail of this distribution beyond $z^{*}$. Mathematically, this is expressed as $\alpha=1-\Phi(z^{*})$. Solving for $z^{*}$, we get: $z^{*}=\Phi^{(-1)}(1-\alpha)$, which is a constant value for different $W$.
For the Aar method, the p-value is calculated by applying a gamma transformation to the sum of correlation values of all the tokens:
@@ 5.1 Theoretical Basis: Gold Index is the Best
We denote $\mu=E[\log(\frac{1}{1-u_{i}})]$ as the expected correction value for a single token. Similar to KGW, we assume $\mu_{1}>\mu$ represents the expected correlation value for the watermarked portion. Due to the complexity of the Gamma-Transformation, the detailed proof can be found in Appendix C. Based on the above proof, we simulated the expected score of the statistic and the corresponding threshold as they vary with $W$ for $L=200$ using real data, as shown in Figure 2. It demonstrates that when the window size corresponds to the gold index, the highest detection rate can be achieved while maintaining an acceptable FPR. Therefore, the watermark dilution effect refers to the situation where the window used for detection is not optimal, resulting in the computed statistics not reaching their peak or minimum, which in turn affects detection performance.
(a) KGW, $\alpha=10^{-6},\gamma=0.5,\gamma_{1}=0.75$
(b) Aar, $\alpha=10^{-6},\mu=1.0,\mu_{1}=1.6$
Figure 2: Expected score of the statistics and the corresponding threshold across various $W$.
5.2 Suspicious Region Localization
Based on the theoretical analysis presented in Section 5.1, the design concept of WaterSeeker employs a coarse-to-fine process to gradually approximate the gold index, thereby achieving high detection performance. In the first step (coarse step), WaterSeeker uses a localization algorithm to identify suspicious watermarked regions. This process narrows the detection target to a small segment that encompasses the ground truth segment, while ensuring relatively small deviation from the gold index. Below is the detailed process of the localization algorithm:
(1) Score List Computation: A small sliding window (i.e., $W=50$) is used to traverse the entire text. For each window position, a statistical measure is computed to reflect the average watermark intensity. In the KGW method, this statistic is $|s|_{G}/W$, representing the density of green tokens within the window. In the Aar method, it is calculated as $\frac{\sum\log(1/(1-u))}{W}$, representing the average correlation value between tokens and the pseudo-random sequence within the window. This results in a score list $s$ of length $N-W+1$, where $s_{i}$ represents the average watermark intensity from $n_{i}$ to $n_{i}+W$.
(2) Anomaly Extraction: In this step, we design an anomaly extraction algorithm to identify outliers from the score list. We calculate the mean of the score list, the top-k mean, and the standard deviation, denoted as $s_{\text{mean}}$, $s_{\text{top-k-mean}}$, and $s_{\text{std}}$, respectively. The extracted outliers are the points that satisfy the following condition, where $\theta_{1}$ and $\theta_{2}$ are hyperparameters:
$\text{score}>s_{\text{mean}}+\max((s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1},s_{\text{std}}\cdot\theta_{2}).$
In samples containing a watermarked segment, the first term in the max expression becomes significant. Introducing the top-k mean allows for adaptive adjustment of the deviation from the mean based on different watermark strengths. Conversely, in non-watermarked samples, the difference between the top-k mean and the mean is relatively small. In this case, to reduce false positives, a point must deviate from the mean by at least $s_{\text{std}}\cdot\theta_{2}$ to be considered an anomaly.
(3) Fragment Connection: In this step, we use a connecting method with a certain tolerance to link nearby outliers, then filter out segments that are too short, returning a list of indices.
Table 1: We compared the detection performance of WaterSeeker with other methods, including Full-text Detection, WinMax (Kirchenbauer et al., 2024), and FLSW. The reported metrics include FPR, FNR, F1 score, and the average IoU between detected segments and ground truth segments in positive samples.
Model Method KGW Aar
FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
Llama-2-7b Full-text Detection 0.000 1.000 [0.987] 0.000 [0.026] 0.000 0.000 1.000 [0.967] 0.000 [0.065] 0.001
WinMax 0.017 0.273 0.834 0.661 0.017 0.410 0.734 0.537
FLSW-100 0.003 0.473 0.688 0.448 0.003 0.547 0.622 0.363
FLSW-200 0.003 0.440 0.716 0.417 0.000 0.577 0.595 0.338
FLSW-300 0.007 0.683 0.479 0.313 0.000 0.783 0.356 0.265
FLSW-400 0.003 0.897 0.187 0.230 0.003 0.933 0.125 0.191
WaterSeeker(Ours) 0.017 0.313 0.806 0.624 0.010 0.440 0.713 0.507
Mistral-7b Full-text Detection 0.000 1.000 [0.990] 0.000 [0.020] 0.000 0.000 1.000 [0.973] 0.000 [0.052] 0.001
WinMax 0.010 0.277 0.835 0.656 0.013 0.370 0.767 0.562
FLSW-100 0.000 0.500 0.667 0.433 0.000 0.497 0.670 0.411
FLSW-200 0.000 0.410 0.742 0.452 0.003 0.560 0.610 0.359
FLSW-300 0.003 0.577 0.593 0.341 0.003 0.827 0.295 0.254
FLSW-400 0.003 0.810 0.318 0.244 0.007 0.943 0.107 0.195
WaterSeeker(Ours) 0.007 0.290 0.827 0.641 0.010 0.390 0.753 0.542
The scores within watermarked segments are relatively stable and close to the top-k mean, while scores outside the watermark remain stable and close to the overall mean due to the dilution effect of large natural text. When using $(s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1}\quad(\theta_{1}<1)$ as a distinction criterion, the extracted abnormal segment’s start and end points ($s^{\prime}$ and $e^{\prime}$) generally satisfy $s^{\prime}\in(s-W,s)$ and $e^{\prime}\in(e,e+W)$. This ensures that the extracted suspicious watermarked regions likely cover the actual segments, with starting and ending deviations within a window size.
5.3 Local Traverse Detection
After obtaining the localization results from the first step, the second phase (fine step), local traverse detection, conducts a more detailed verification. For each $(s^{\prime},e^{\prime})$ pair in the localization results, traverse inwards through the segments where the start point falls within $[s^{\prime},s^{\prime}+W)$ and the end point within $(e^{\prime}-W,e^{\prime}]$. Based on the previous analysis of the ranges of $s^{\prime}$ and $e^{\prime}$, this traversal has a high probability of reaching the gold index. Perform full-text detection sequentially on these segments, and select the statistic with the most significance to compare with the threshold. The pseudocode for the entire WaterSeeker algorithm can be found in Algorithm 3.
5.4 Time Complexity Analysis
Time Complexity of WaterSeeker: Suspicious Region Localization has a time complexity of $O(N)$. This step involves score list computation, anomaly detection, and fragment connection, each with a time complexity of $O(N)$, where $N$ is the text length. Local Traverse Detection has a time complexity of $O(W^{2})$. For each suspicious region, this step examines $W^{2}$ windows, where $W$ is the window size. The total time complexity of WaterSeeker is $O(N+W^{2})$. In practice, $W^{2}$ is typically kept lower than $N$, as a slightly larger window (i.e., $W=50$, detailed in Appendix E) suffices for a smooth and low-noise representation of the surrounding watermark intensity. Thus, the overall time complexity of WaterSeeker is $O(N)$.
Lower Bound Complexity for the Problem: To detect watermarked segments in a long text, any algorithm must examine each token in the text at least once. This requirement establishes a lower bound of $\Omega(N)$ for the time complexity of the problem, as at least one full pass through the text is necessary. Consequently, the WaterSeeker algorithm achieves a time complexity that matches the theoretical lower bound of the problem.
6 Experiment
(a) KGW, $\delta=2.0$
(b) KGW, $\delta=1.5$
(c) KGW, $\delta=1.0$
(d) Aar, temp=0.3
(e) Aar, temp=0.2
(f) Aar, temp=0.1
Figure 3: This figure compares the detection performance of WaterSeeker and FLSW under varying watermark lengths and strengths, using Llama-2-7b as the generation model.
Table 2: Average detection time per sample for various detection methods ($N\approx 10,000$, Unit: s).
Full-text Detection WinMax FLSW-100 FLSW-200 FLSW-300 FLSW-400 WaterSeeker
KGW 1.70 14.16 1.76 1.76 1.79 1.80 1.75
Aar 0.54 2733.78 1.62 1.65 1.66 1.64 1.68
6.1 Experiment Settings
Watermarking Methods and Language Models: We selected two representative watermarking algorithms, KGW (Kirchenbauer et al., 2023) and Aar (Aaronson and Kirchner, 2022), each at three strength levels. KGW’s strength was set by the $\delta$ parameter (2.0=strong, 1.5=medium, 1.0=weak), while Aar’s strength used the temperature parameter (0.3=strong, 0.2=medium, 0.1=weak). We used Llama-2-7b (Touvron et al., 2023) and Mistral-7b (Jiang et al., 2024) as generation models.
Dataset Construction: The first 30 tokens of each entry in the C4 dataset (Raffel et al., 2020) were used for prompts. Watermarked segments of random length (100 to 400 tokens) were then generated using randomly selected watermark strengths. For positive examples, one such segment was randomly inserted into each 10,000-token Wikipedia passage (Foundation, ). Negative examples consist of unmodified 10,000-token Wikipedia corpus. Based on this procedure, four datasets were created, each containing 300 positive and 300 negative examples: KGW-llama, KGW-mistral, Aar-llama, and Aar-mistral. The specific distributions of watermark strengths and lengths are detailed in Appendix F.
Baselines: As introduced in Section 4, Full-text Detection and WinMax (Kirchenbauer et al., 2024) are chosen, as well as Fix-Length Sliding Window method with $W$ of 100, 200, 300 and 400.
Hyper-parameters: The parameters related to WaterSeeker are as follows: $W=50,k=20$, with a tolerance for fragment connection set to 100. The parameter $\theta_{1}$ is set to 0.5, and $\theta_{2}$ is set to 1.5 for both algorithms. The threshold selection within the specified window is detailed in Appendix D. Notably, careful threshold selection is crucial for maintaining an acceptable false positive rate, as traversing long texts is prone to accumulating false positives. Evaluation-related parameters: $\theta$ mentioned in Section 3.2 is set to 0.5.
Table 3: This table presents the contributions of the first step of WaterSeeker: Suspicious Segment Localization. It lists the average coverage of localization results relative to the ground truth segments under various watermark algorithms and watermark strengths, as well as the average offsets of the detected start and end indices.
Metrics KGW Aar
$\delta=2.0$ $\delta=1.5$ $\delta=1.0$ temp=0.3 temp=0.2 temp=0.1
Average Coverage 0.992 0.984 0.953 0.996 0.991 0.977
Average Offset (Start) 20.799 30.057 13.992 22.559 30.546 20.624
Average Offset (End) 25.985 25.299 28.030 29.459 33.059 30.304
@@ 6.1 Experiment Settings
Table 4: Comparison of detection performance with and without Local Traverse Detection across four datasets.
Model Setting KGW Aar
FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
Llama-2-7b w. local traverse 0.017 0.313 0.806 0.624 0.010 0.440 0.713 0.507
w/o local traverse 0.003 0.390 0.756 0.520 0.000 0.450 0.710 0.462
Mistral-7b w. local traverse 0.007 0.290 0.827 0.641 0.010 0.390 0.753 0.542
w/o local traverse 0.003 0.380 0.763 0.534 0.003 0.413 0.738 0.490
6.2 Detection Performance Analysis
In Table 1, we present the detection performance of WaterSeeker alongside various baseline algorithms, evaluated on four datasets. Given that full-text detection lacks localization capabilities, its FNR is necessarily 1 under the evaluation rule requiring IoU $>\theta$ for successful detection. Therefore, we additionally report the FNR and corresponding F1 score for Full-text Detection without the IoU $>\theta$ constraint, highlighted in red. Additional performance results for WaterSeeker under various mixing ratios are presented in Appendix G.
The results demonstrate that WaterSeeker achieves lower FNR and higher F1 score, significantly surpassing other watermark detection algorithms, including Full-text Detection, FLSW-100, FLSW-200, FLSW-300, and FLSW-400, while being comparable to WinMax. Based on the principle that WinMax is guaranteed to reach the gold index by evaluating all possible windows, it represents the upper bound of detection performance. However, subsequent experiments show that WinMax has extremely high time complexity (Section 6.4). Furthermore, the IoU results in the table reveal that WaterSeeker possesses a significantly higher localization capability compared to the baselines, again comparable to WinMax.
6.3 Adaptability Analysis
While the main experiment used four datasets with a mix of watermarked segments of varying strengths and lengths, this section compares the detection capabilities of WaterSeeker and the FLSW method across specific watermark strengths and lengths. Figure 3 shows that the performance of different detection algorithms follows a consistent trend: watermarked segments with higher strengths and longer lengths are more easily detected. FLSW performs well within a length range close to its window size but shows poor adaptability in other ranges due to its fixed-length nature, which leads to dilution effects when detecting longer or shorter watermarked segments. In contrast, WaterSeeker exhibits good adaptability across different length ranges by using anomaly extraction techniques and incorporates a top-k score mechanism to adapt to varying watermark strengths.
6.4 Time Complexity Analysis
Table 2 shows the average time spent per sample for WaterSeeker and other baseline algorithms during detection. WaterSeeker’s time is comparable to full-text detection and FLSW, and significantly lower than WinMax. Sections 4.2 and 5.4 analyze the theoretical time complexity of WinMax and WaterSeeker, respectively. This section offers a more comprehensive analysis based on experimental results, considering constant factors and investigating the significant disparity in WinMax’s execution time between KGW and Aar.
Let $t_{1}$ be the time to compute a score in the score list, and $t_{2}$ be the time to compute the watermark detection statistic within a local window. WinMax’s time expenditure to detect a sample is $N\cdot t_{1}+N\cdot(W_{\max}-W_{\min})\cdot t_{2}$. For WaterSeeker, the localization step consists of score list computation, anomaly detection, and fragment connection, each with complexity $O(N)$. Denoting the time for the latter two sub-steps as $t_{3}\cdot N$, the total time for the first step is $N\cdot t_{1}+N\cdot t_{3}$. The local traverse detection step requires $W^{2}\cdot t_{2}$ time. Thus, WaterSeeker’s total time expenditure is $N\cdot t_{1}+N\cdot t_{3}+W^{2}\cdot t_{2}$.
For the Aar (Aaronson and Kirchner, 2022) watermarking method, calculating the watermark detection statistic involves a complex GammaTransform function, resulting in a larger $t_{2}$. This leads to a significant time difference between WinMax and WaterSeeker due to the large disparity in coefficients preceding $t_{2}$. In this experiment, with $N=10,000,W_{\max}=400,W_{\min}=100,W^{2}=2,500$, the coefficient difference is a factor of 1,200. For the KGW (Kirchenbauer et al., 2023) watermarking method, calculating the z-score within a local window is relatively straightforward, reducing the time difference between WinMax and WaterSeeker.
WaterSeeker is a general watermark detection method with consistently low time expenditure across various watermarking algorithms. In contrast, WinMax exhibits extremely high time costs when calculating the watermark detection statistic is complex, making it challenging for practical use.
6.5 Ablation Study
@@
We analyze the effectiveness of the two stages of WaterSeeker through an ablation study. The first stage, Suspicious Segment Localization, aims to achieve high coverage of the ground truth segments while keeping the start and end offsets within a specified window size. This ensures subsequent local traversals can reach the gold index. Table 3 shows that Step 1 achieves an average coverage exceeding 0.95 across various watermark algorithms and strengths, with average start and end offsets less than 50, remaining within the designated window size. Step 1 fulfills its purpose effectively.
@@
Local Traverse Detection performs a localized traversal based on the segments from Step 1, allowing for more refined verification within the window. Table 4 shows that across different LLMs and watermarking algorithms, Local Traverse consistently enhances detection F1 score and average IoU compared to directly applying full-text detection with the localization results, making it an indispensable component of WaterSeeker.
@@ 7 Conclusion
While our method has demonstrated effectiveness in detecting watermarked segments within large documents, there are still some limitations that need to be addressed in future work. First, from an evaluation perspective, due to resource constraints, we only conducted experiments on Llama-2-7B and Mistral-7B models. The effectiveness of our method on larger and more powerful models remains to be further verified. Second, WaterSeeker’s performance may decrease with very short or weak watermarks. Enhancing the sensitivity of WaterSeeker to detect shorter and weaker watermarks is an area for future improvement, which may involve refining the anomaly extraction algorithms or incorporating additional contextual analysis. Lastly, parameter selection, such as threshold settings, is crucial and can be challenging in different environments. Currently, parameters are manually tuned based on observations, so developing adaptive tuning methods, potentially using machine learning, would enhance WaterSeeker’s robustness and practicality across diverse scenarios.
@@ A.2 Aar
Appendix B Pseudocode of WaterSeeker and Detection Baselines
Pseudocode of WaterSeeker, WinMax and FLSW could be found in Algorithm 3, 1 and 2, respectively.

### v4

@@ 6.5 Ablation Study
[…]
@@
Table 6: Comparison of detection performance with and without Local Traverse Detection.
Settings KGW Aar
F1 IoU F1 IoU
w. Traversal 0.872 0.675 0.819 0.578
w/o Traversal 0.817 0.576 0.765 0.509
We analyze the effectiveness of the two stages of WaterSeeker through an ablation study.
@@
Stage 1: Suspicious Region Localization. This stage aims to narrow down the detection target from a large document to a smaller region. The goal is to achieve high coverage of the ground truth segments while maintaining the start and end offsets within a specified window size. This ensures that subsequent local traversals can access the gold index. As shown in Table 5, Step 1 achieves an average coverage exceeding 0.9, with average start and end offsets remaining below $W$ across various watermark algorithms and strengths, demonstrating good adaptability.
@@
Stage 2: Local Traverse Detection. Local Traverse Detection performs a localized iteration based on the segments from Stage 1, allowing for more refined verification within the window. Table 6 shows that across different watermarking algorithms, Local Traverse consistently enhances detection F1 score and average IoU compared to directly applying detection with the localization results, making it an indispensable component of WaterSeeker.
6.6 Robustness against Text Edit Attacks
Figure 4: Robustness of WinMax, FLSW, and WaterSeeker against text edit attacks. The detection F1 score is reported for no attack, random word deletion attack (ratio = 0.3), and random word substitution attack (ratio = 0.3), along with the time cost for processing one sample.
In this section, we further examine WaterSeeker’s robustness against text edit attacks, as watermarked segments generated by LLMs may be altered before integration into non-watermarked documents.
Figure 4 illustrates the detection robustness of WaterSeeker, WinMax and FLSW against random word deletion (ratio=0.3) and substitution attacks (ratio=0.3, utilizing WordNet (Fellbaum, 1998) for synonym sets). The figure also contrasts these results with those obtained under no attack, along with the associated time costs. In this experiment, the KGW watermarking algorithm was employed, and Llama-2-7b is utilized as generation model. Since text editing can weaken the strength of the watermark, the watermark fragment intensity was set to a strong level under the "No attack" condition, specifically $\delta=2.0$.
It can be observed from the figure that WaterSeeker achieves strong robustness against word deletion and word substitution attacks, with an F1 score exceeding 0.9. Compared with other baselines, only WinMax-1 perform slightly better than WaterSeeker, but at a significant cost in terms of time.
@@ 7 Conclusion
While our method has demonstrated effectiveness in detecting watermarked segments within large documents, there are still some limitations that need to be addressed in future work. First, from an evaluation perspective, due to resource constraints, we only conducted experiments on Llama-2-7B and Mistral-7B models. The effectiveness of our method on larger and more powerful models remains to be further verified. Second, WaterSeeker’s performance may decrease with very short or weak watermarks. Enhancing the sensitivity of WaterSeeker to detect shorter and weaker watermarks is an area for future improvement, which may involve refining the anomaly extraction algorithms or incorporating additional contextual analysis. Lastly, parameter selection, including threshold settings to control false positives, is crucial and can be challenging in different environments. Stricter threshold controls can reduce the detection rate, necessitating adjustments based on the specific requirements of the actual settings.
@@ A.2 Aar
Appendix B Supplementary Experimental Results Using Mistral-7b
Table 7: We evaluated the detection performance of WaterSeeker against various methods, including Full-text Detection, WinMax (Kirchenbauer et al., 2024), and FLSW. Results in this table utilized Mistral-7b as the generation model. The metrics reported include false positive rate (FPR), false negative rate (FNR), F1 score, average Intersection over Union (IoU) between detected and ground truth segments, and processing time per sample. Best performances are highlighted in bold, while the second-best are underlined.
Method KGW Aar
[…]
WinMax-1 0.010 0.243 0.857 0.641 1632.11 0.013 0.297 0.819 0.588 3615.42
WinMax-50 0.007 0.270 0.841 0.613 34.31 0.007 0.320 0.806 0.560 72.13
WinMax-100 0.003 0.283 0.833 0.588 17.16 0.007 0.333 0.797 0.531 35.34
WinMax-200 0.003 0.340 0.794 0.511 9.12 0.003 0.347 0.789 0.501 18.38
WaterSeeker 0.010 0.253 0.850 0.634 1.75 0.010 0.300 0.819 0.563 0.41
FLSW-100 0.000 0.463 0.698 0.393 1.76 0.000 0.473 0.690 0.367 1.31
FLSW-200 0.000 0.373 0.770 0.426 1.76 0.003 0.387 0.759 0.412 1.29
FLSW-300 0.003 0.437 0.719 0.331 1.76 0.003 0.387 0.759 0.334 1.29
FLSW-400 0.003 0.540 0.629 0.218 1.75 0.007 0.440 0.715 0.249 1.27
Supplementary results for main experiment using Mistral-7b as generation model is shown in Table 7, showing consistent trend with Llama-2-7b.
Appendix C Pseudocode of Detection Baselines
Pseudocode of WinMax and FLSW could be found in Algorithm 2 and 3, respectively.